40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3856     Entry time: Wed Nov 3 19:14:00 2010
Author: rana 
Type: Update 
Category: PSL 
Subject: PMC mode matching update 

I moved the lens just before the PMC to check the mode matching landscape. The PMC trans went up from ~6.5 to ~6.8. That's 5% with ~1 hour of work.

As per the micrometer, this took ~7-8 mm of travel. Since there's so much power left in the HOMs, we we we will have to do a proper mode scan and re-calculate the solution.

The measured transmission is now ~610 mW. The power reflected from the PMC with it unlocked is ~1400 mW.

ELOG V3.1.3-